# MGAS: Multi-Granularity Architecture Search for Trade-Off Between Model Effectiveness and Efficiency

### v3

@@ V-D2 Results on ImageNet
Similarly, we employ two different settings for ImageNet. To compare with coarse-grained methods, we explore large-size models with 14 cells and an initial filter number of 48. To compare with fine-grained methods, given their high search cost, we directly evaluate the small-size models discovered on CIFAR-10 using both MGAS and the baseline DASS instead of searching from scratch on Imagenet. In this way, we can also assess the generalization ability of the discovered architecture. We retrain the architectures for 250 epochs for evaluation, and the results are provided in Table II.
@@
As shown in the table, MGAS surpasses all the baselines in terms of top-1 accuracy density, which indicates that MGAS can be generalized well to large datasets. Moreover, the transfered architecture discovered by MGAS demonstrates remarkably higher accuracy compared to that discovered by DASS. This highlights the strong generalization capability of the architectures generated by our method.
@@ V-E Adaptive Pruning Consistently Achieves a Better Granularity Balance at Different Model Sizes (RQ2)
The superiority of MGAS over methods with manually defined ratios has already been demonstrated for certain model sizes. In this section, we verify whether adaptive pruning can consistently achieve a better granularity balance for different target model sizes.
To simplify the process, we only conduct operation-level and filter-level search. To eliminate the effect of multi-stage search, we employ the small search space setting with 8 cells and an initial filter number of 24, where the constructed super-net is small enough to be directly put into the memory without decomposition. Through one search process, we generate a series of architectures of varying sizes using adaptive pruning. Afterwards, we calculate the average pruning rates at both the operation and filter levels. The average pruning rates observed are 36 operations and 0.7% filters per two epochs. Based on this, we implement two basic gradual pruning methods with fixed pruning rates as baselines. Baseline 1 has higher operation-level pruning rates and lower filter-level pruning rates, while baseline 2 has higher filter-level pruning rates and lower operation-level pruning rates:
@@
Employing these baseline methods, we obtain baseline architectures that exhibit variations in the remaining ratios at operation and filter levels. Here, the remaining ratio is defined as the ratio of the remaining unit number to the original unit number. For operations, the original unit number refers to the initial operation number in the super-net. For filters, the original unit number represents the initial filter number in operations within the discovered architecture.
The remaining ratios in the discovered architectures of adaptive pruning and the baselines are depicted in Fig. 5. As expected, baseline 1 generally retains more filters and fewer operations compared to adaptive pruning at similar model sizes. Conversely, baseline 2 preserves more operations and fewer filters. Moreover, it can be observed that adaptive pruning generates slimmer architectures at sizes of 1.05M and 1.17M, while producing wider architectures at smaller model sizes. This illustrates its ability to adaptively adjust the relative remaining ratios of operations and filters as the architecture evolves.
The performance of these methods is illustrated in Fig. 6. As displayed in the figure, adaptive pruning outperforms both baselines at all the model sizes, indicating its ability to consistently achieve a superior balance between granularities.
Figure 5: Remaining ratios of operations and filters in pruning with fixed pruning rates and our proposed adaptive pruning.
Figure 6: Performance of pruning with fixed pruning rates and our proposed adaptive pruning.
@@ V-F1 Reduction of memory consumption
To evaluate the reduction in memory consumption, we conduct a comparison between the search processes using both single-stage and multi-stage settings on CIFAR-10 in the small search space. The results are presented in Table III.
As shown in the table, the memory consumption reaches its peak at stage 3 in the multi-stage search, which is only 54% of the memory consumption observed in the single-stage search.
V-F2 Effect of hyper-parameter $r$
To assess the significance of the regrowing process and the effect of the hyper-parameter $r$, we conduct experiments comparing various regrowing ratios on CIFAR-10 in the large search space. The results are presented in Table IV.
From the table, it is apparent that the absence of the regrowing process, represented by $r$ equaling 0, leads to limited performance. This is reasonable since units that have been pre-maturely pruned cannot be recovered. As the value of $r$ increases, there is a corresponding improvement in accuracy from 97.20% to 97.34%. Such improvement can be attributed to the greater number of pruned units being re-evaluated in subsequent stages.
@@ V-F2 Effect of hyper-parameter $r$
V-F3 Preservation of search quality
To assess whether multi-stage search can consistently preserve high search quality, we compare it with other memory reduction strategies in different variations of the search space. To enable the comparison with proxy strategies, only operation-level search is conducted. To avoid any other confounding factors, adaptive pruning is not applied.
We choose DARTSV1, PC-DARTS and GDAS as baseline methods to represent depth-level proxy, channel-level proxy and decomposition strategies, respectively. We evaluate these methods on three 8-cell search spaces with an initial filter number of 36. Search space 1 represents the original DARTS search space. Search space 2 relaxes the connection constraint, allowing for the discovery of multi-path architectures. Search space 3 further relaxes the identical cell constraint, enabling different cell structures. The results are shown in Table V.
In search space 1, multi-stage search outperforms both DARTSV1 and GDAS in terms of accuracy and achieves similar performance with PC-DARTS, highlighting its effectiveness in preserving high search quality. After removing the manual constraints on the search spaces, the accuracy of the architectures discovered by both proxy and decomposition methods decrease significantly. Specifically, the path-level decomposition method GDAS performs poorly in Search Space 2, indicating its limited ability to explore multi-path architectures. Similarly, PC-DARTS, which performs well in search space 1, experiences a significant accuracy drop in search spaces 2 and 3, revealing the proxy strategy’s failure in preserving search quality when the search space exhibits more flexibility. In comparison, multi-stage search maintains high search quality in all the three search spaces.
@@ V-F3 Preservation of search quality
Figure 7: Distribution patterns of local units in discovered architectures. This figure illustrates the variance in unit distribution across different cells and operation types within the neural network architectures discovered on CIFAR-10 and ImageNet datasets. Notably, it highlights the preferences for fewer but wider and denser convolution operations in reduction cells, the predominance of $5\times 5$ separable convolution operation, and the variance in operation width and density between the middle and ends of large architectures. Differences in pattern preferences between the CIFAR-10 and ImageNet datasets are also depicted, reflecting the adaptation of the architecture to dataset complexity.
V-G Analysis of the Local Patterns in Discovered Architectures
@@
In this section, we investigate the local unit distribution regarding the convolution operations inside the discovered architectures from Table I and II, aiming to provide insights for neural network design. The remaining ratios of units at different granularity levels regarding different cells and operations in the discovered architectures are demonstrated in Fig. 7.
Upon examination of the three architectures, it can be observed that reduction cells (cell 3 and 6 in the small architecture, cell 5 and 10 in the large architectures) tend to preserve wider and denser convolution operations, while the operation number is not necessarily large compared to other cells. In terms of operation types, separable convolution with a kernel size of $5\times 5$ generally appears to be more effective.
The pattern variance between different cells and operation types is relatively small in the small architecture, whereas it is more pronounced in the large architectures. In the large architectures, the middle cells require higher operation density rather than width, while both ends of the architectures favor width over density. These common characteristics reasonably explain the generalization ability exhibited by the architecture discovered on CIFAR-10 to ImageNet.
There are still slight differences in patterns between the architectures discovered on CIFAR-10 and ImageNet. While different operations display similar patterns on the smaller CIFAR-10 dataset, they exhibit distinct preferences for the more complex ImageNet dataset. When necessary, reducing filters of separable convolution operations while maintaining density may have minimal adverse effects. In contrast, dilated convolution operations can tolerate higher levels of sparsity. These findings align with their inherent functionalities, as separable convolution emphasizes preservation of fine-grained local features with its smaller receptive field, while dilated convolution aims to capture global features from multiple perspectives using a larger receptive field, where intricate details are of less importance.
@@ VI Conclusion
In this work, we introduce a new perspective to discover effective yet parameter-efficient neural networks by comprehensively exploring searchable units of multiple granularities and optimizing the relative numbers of the remaining units at different granularity levels. Additionally, we effectively tackle the challenge of high memory consumption of the search process without compromising the search quality via super-net decomposition with progressive re-evaluation. Experimental results demonstrate the consistent superiority of MGAS over other alternatives in terms of the accuracy-parameter trade-off. However, it is important to note that MGAS suffers from a lengthy search time due to the time-consuming convergence of the super-net. To address this limitation, our future research efforts will focus on integrating zero-cost metrics to estimate the importance of searchable units at initialization.

### v4

@@ V-D2 Results on ImageNet in DARTS search space
Similarly, we employ two different settings for ImageNet. For the large model setting, we explore models with 14 cells and an initial filter number of 48. For the small model setting, given the high search cost of the baseline DASS, we directly transfer the small-size architectures discovered on CIFAR-10. The results are provided in Table II.
@@
As shown in the table, MG-DARTS surpasses all the baselines in terms of accuracy density, which indicates that MG-DARTS can be generalized well to large datasets. The transfered architecture discovered by MG-DARTS also demonstrates higher accuracy compared to that discovered by DASS. This highlights the strong generalization capability of the architectures generated by our method.
Additionally, following GOLD-NAS [40], our method adopts the basic one-level relaxation of the DARTS optimization strategy without incorporating any regularization techniques. This design choice was made to ensure simplicity and clearly highlight the core contributions of our work. It is worth noting that many well-established regularization techniques, such as $\beta$-DARTS [50], are orthogonal to our framework and can be seamlessly integrated to further improve performance.
V-D3 Results on ImageNet in MobileNet-like search space
To evaluate the effectiveness of our method across different search spaces, we also implement it in a MobileNet-like search space. As shown in Table III, MG-DARTS achieves comparable top-1 accuracy on ImageNet while reducing model size by incorporating weight-level granularity and determining appropriate retention ratios through joint discretization.
@@ V-E Adaptive Pruning Consistently Achieves a Better Granularity Balance at Different Model Sizes (RQ2)
The superiority of MG-DARTS over methods with manually defined ratios has already been demonstrated for certain model sizes. In this section, we verify whether adaptive pruning can consistently achieve a better granularity balance for different target model sizes. To eliminate the effect of multi-stage search, we employ a smaller search space setting with 8 cells, where the constructed super-net is small enough to be directly put into the memory without decomposition.
In this section, the retention ratio is defined as the ratio of the remaining unit number to the original unit number. For operations, the original unit number refers to the initial operation number in the super-net. For filters, the original unit number denotes the initial filter number in operations within the discovered architecture. For weights, the original unit number denotes the initial weight number in filters within the discovered architecture.
V-E1 Comparison with different retention ratios
In this section, we conduct the search using combinations of any two granularity levels and manually vary the retention ratios of units at each level to evaluate whether our adaptive pruning method effectively determines appropriate ratios. The results are shown in Fig. 5, where the ratios discovered by our method are marked with a star.
For the operation-filter combination, we increase the filter retention ratio by 10% at each step and adjust the operation count accordingly to maintain a similar overall model size. Similarly, for the operation–weight and filter–weight combinations, we increase the weight retention ratio by 10% and 5% respectively, while adjusting the corresponding filter or operation count. This setup allows us to verify whether the discovered ratios fall within high-performance regions.
Note that the number of filters found during the search may slightly differ from those in the final architecture due to post-search filter alignment. To ensure consistency, we report the retention ratios obtained directly from the search phase. While the exact numerical ratios may not always match, the relative proportions of filters across different settings remain consistent, ensuring a fair comparison.
As shown in the figure, the optimal retention ratio varies across different granularity combinations, with performance peaks appearing at several points. Nevertheless, adaptive pruning effectively identifies configurations that lie close to these peaks and yields top-tier performance across all settings.
Figure 5: Comparison with different manual ratios. Adaptive pruning effectively identifies configurations that lie within high-performance region.
V-E2 Comparison with different pruning strategies
To simplify the process, we only conduct search at the operation and filter levels. We compare our method with the widely used progressive pruning strategy that relies on predefined pruning rates. Through a single search process, our method generates a series of architectures with varying sizes using adaptive pruning. We then calculate the average pruning rates at both the operation and filter levels, which are approximately 36 operations and 0.7% filters pruned every two epochs.
Based on these observed rates, we design two baseline methods to serve as comparisons. This ensures that the overall retention ratios of the baselines are reasonably close to our method, so that our advantages can be better highlighted. Specifically, baseline 1 adopts a higher pruning rate at the operation level and a lower rate at the filter level, while baseline 2 adopts a higher filter-level pruning rate and a lower operation-level pruning rate:
@@ V-E2 Comparison with different pruning strategies
Employing these baseline methods, we obtain baseline architectures that exhibit variations in the retention ratios at operation and filter levels.
The retention ratios in the discovered architectures of adaptive pruning and the baselines are depicted in Fig. 6. As expected, baseline 1 generally retains more filters and fewer operations compared to adaptive pruning at similar model sizes. Conversely, baseline 2 preserves more operations and fewer filters. Moreover, it can be observed that adaptive pruning generates slimmer architectures at sizes of 1.05M and 1.17M, while producing wider architectures at smaller model sizes. This illustrates its ability to adaptively adjust the retention ratios of operations and filters as the architecture evolves.
The performance of these methods is illustrated in Fig. 7. As shown in the figure, adaptive pruning consistently outperforms both baselines across all model sizes, forming a Pareto frontier. This indicates its effectiveness in dynamically determining appropriate retention ratios for different target complexities.
Figure 6: Retention ratios of operations and filters in pruning with fixed pruning rates and our proposed adaptive pruning.
Figure 7: Comparison with progressive pruning with fixed pruning rates. Adaptive pruning consistently yields models with improved performance.
V-E3 Ablation of different granularity levels
To evaluate the importance of each granularity in our method, we conduct experiments comparing architectures searched with different combinations of granularity levels. As shown in Table VIII, incorporating filter-level and weight-level searchable units leads to improved model performance at similar model size, as redundancy is better identified and reduced. The best performance is achieved when all three levels are jointly considered.
TABLE VIII: Performance Comparison of search spaces with different granularity levels.
Search Space Accuracy (%) Params (M)
operation 93.11 0.7
operation + filter 93.21 0.71
operation + weight 93.37 0.72
operation + filter + weight 93.52 0.71
TABLE IX: Performance Comparison of different optimization strategies.
   Strategy    Accuracy (%)    Params (M)
   Iterative (Full Epoch)    97.13    2
   Iterative (Half Epoch)    97.19    2.2
   Joint    97.47    2.2
V-E4 Effect of the optimization strategy
To optimize the granularity parameters $g$ (i.e., $\alpha,\beta,\omega$) and their associated thresholds $t_{g}$ (i.e., $t_{\alpha},t_{\beta},t_{\omega}$), a natural idea is to iteratively optimize $g$ and $t_{g}$ in an alternating manner to avoid conflicting gradient signals. However, as shown in Table IX, we empirically find that the joint optimization strategy, where both $g$ and $t_{g}$ are updated simultaneously, achieves the best performance. It obtains the highest accuracy of 97.47%, outperforming the iterative strategies with half-epoch (97.19%) and full-epoch (97.13%) update intervals.
We attribute the inferior performance of iterative strategies to two main issues: error propagation and reduced update frequency. Specifically, iterative updates are designed to approximate the following bi-level optimization problem:
$\min_{t_{g}}\ \mathcal{L}(g^{*}(t_{g}),t_{g}),$ (11)
$\text{where}\quad g^{*}(t_{g})=\arg\min_{g}\ \mathcal{L}(g,t_{g})$
Since $g^{*}(t_{g})$ cannot be solved exactly, iterative methods rely on partial updates of $g$, which results in inaccurate gradients for $t_{g}$ [39], leading to error propagation. Moreover, as only one parameter group is updated at a time, both $g$ and $t_{g}$ receive only half as many updates as they would under joint optimization under the same training budget, which slows down convergence and degrades the performance.
@@ V-F1 Reduction of memory consumption
To evaluate the reduction in memory consumption, we conduct a comparison between the search processes using both single-stage and multi-stage settings on CIFAR-10 under the small search space configuration. The results are presented in Table IV.
As shown in the table, the peak memory consumption in the multi-stage search (12.4G) is only 46.9% of that in the single-stage counterpart (26.4G), with merely 2 additional GPU hours incurred and no accuracy drop observed. This indicates that the multi-stage design effectively reduces memory usage at a modest computational cost, without compromising performance.
V-F2 Preservation of search quality
To assess whether our multi-stage search can consistently preserve high search quality, we compare it with other memory reduction strategies in different variations of the search space. For simplicity, only operation-level search is conducted.
We choose DARTSV1, PC-DARTS and GDAS as baseline methods to represent depth-level, width-level and path-level strategies, respectively. We evaluate these methods on three 8-cell search spaces with an initial filter number of 36. Search space 1 represents the original DARTS search space. Search space 2 relaxes the connection constraint, allowing for the discovery of multi-path architectures. Search space 3 further relaxes the identical cell constraint, enabling different cell structures. The results are shown in Table VI.
In search space 1, multi-stage search outperforms both DARTSV1 and GDAS in terms of accuracy and achieves similar performance with PC-DARTS, highlighting its effectiveness in preserving high search quality. After removing the manual constraints on the search spaces, the accuracy of the architectures discovered by baseline methods decrease significantly. Specifically, the path-level method GDAS performs poorly in Search Space 2, indicating its limited ability to explore multi-path architectures. Similarly, PC-DARTS, which performs well in search space 1, experiences an obvious accuracy drop in search spaces 2 and 3, revealing the width-level strategy’s failure in preserving search quality when the search space exhibits more flexibility. In comparison, our multi-stage search maintains high search quality in all the three search spaces.
V-F3 Effect of hyper-parameter $r$
To assess the significance of the regrowing process and the effect of the hyper-parameter $r$, we conduct experiments comparing various regrowing ratios on CIFAR-10 in the large search space. The results are presented in Table V.
From the table, it is apparent that the absence of the regrowing process, represented by $r$ equaling 0, leads to limited performance. This is reasonable since units that have been pre-maturely pruned cannot be recovered. As the value of $r$ increases, there is a corresponding improvement in accuracy from 97.14% to 97.47%.
@@ V-F3 Effect of hyper-parameter $r$
V-F4 Effect of intermediate task-specific auxiliary layers
To investigate the impact of intermediate auxiliary layers in the multi-stage search strategy, we compare three different configurations, all of which primarily consist of two convolutional layers followed by a dense layer as the classifier. Our design principle is to align the feature map size with the final output while keeping the auxiliary layers lightweight, ensuring that the primary focus of optimization remains on the search space. Their size is kept around 10% of the main part of the search space, ensuring that the total parameter count, including both the discovered architecture at the current stage and its following auxiliary layers, remains approximately the same as that of the final discovered architecture.
We denote the configuration used in our method as the medium size. Based on this, we create two variants: the large size, where we increase the kernel size of the convolutional layers, and the small size, where we reduce the number of channels in the convolutional layers. To evaluate the performance under these different configurations, we train the discovered architectures for 200 epochs.
As shown in Table X, the auxiliary task-specific layers need to be within a reasonable size range to ensure good performance. Their primary purpose is to simulate high-level feature extraction and prediction for the earlier stages, allowing the early-stage layers to focus on capturing low-level features. If these auxiliary layers are too small, the final layers of the early-stage sub-nets may be forced to evolve into high-level feature extractors themselves, which disrupts the intended division of labor and degrades overall performance.
TABLE X: Performance comparison of different auxiliary layers.
  Component   Params (M)   Accuracy (%)
  Small Aux   0.69   94.26
  Medium Aux   2.69   94.67
  Large Aux   3.23   94.76
  Main Parts   21   -
Figure 8: Dynamics of thresholds and retention ratios at different granularity levels during search.
@@ V-F4 Effect of intermediate task-specific auxiliary layers
Figure 9: Distribution patterns of local units in discovered architectures. This figure illustrates the variance in unit distribution across different cells and operation types within the neural network architectures discovered on CIFAR-10 and ImageNet datasets. Notably, it highlights a tendency to preserve more operations in the early layers and more filters at both ends, as well as a general preference for separable over dilated convolution operations.
V-G Visualization
To further understand the proposed MG-DARTS search process and the resulting architectures, we provide two visual analysis. The first part illustrates the evolution of the architectures throughout the search process, while the second part presents the architectural patterns of the discovered architectures.
V-G1 Architectural Evolution during Search
Fig. 8 visualizes the epoch-wise evolution of architectural parameters during the search process under the small search space configuration. The upper plot depicts the progression of the average learned thresholds, while the lower plot presents the corresponding retention ratios.
The thresholds for operation and filter levels exhibit a relatively smooth and gradual variation, whereas the weight-level threshold remains low for early epochs and rises sharply later, indicating delayed pruning at the weight level. From the retention ratio curves, we can observe a clearer coarse-to-fine pruning pattern, where the operation and filter ratios decrease quickly during the early epochs, while the weight ratio stays close to 1.0 initially and only begins to decline significantly in the later phase.
This behavior can be attributed to the inherent stability differences among parameters at different granularity levels. In the early phase, the super-net is under-optimized and subject to noisy gradient signals. Coarser-grained structures, due to their limited candidate space and larger influence on performance, tend to exhibit more consistent and distinguishable importance even under such unstable conditions. In contrast, finer-grained structures are more sensitive to noise, causing their thresholds to receive less consistent gradients and update more slowly. This inherent coarse-to-fine pruning pattern naturally contributes to a more stable and reliable search process.
V-G2 Architectural Patterns in Discovered Architectures
@@
In this section, we investigate the local unit distribution of convolution operations within the discovered architectures from Table I and II to uncover general structural patterns. Figure 9 presents the retention ratios of units at different granularity levels, organized by layer depth and operation type.
At the layer level, more operations are kept in the shallower part of the network, which is reasonable since these regions are responsible for processing raw input features and benefit from diverse transformation pathways to extract low-level information. At the filter level, a U-shaped distribution is observed: fewer filters are maintained in the middle, while both ends retain more. This pattern is consistent with previous pruning studies [53], which suggest that filter redundancy tends to be higher in the middle part of the network. Correspondingly, more weights are preserved in the middle, partially due to the reduced number of filters, resulting in denser parameter allocation per filter to maintain sufficient expressiveness.
At the operation-type level, separable convolutions are consistently preferred over dilated ones, in line with findings from most prior differentiable architecture search works [1, 30]. This preference is also reflected in the greater allocation of filters and weights to separable convolutions, highlighting their central role in the resulting architectures.
@@ VI Conclusion
In this work, we introduce a new perspective to discover effective yet parameter-efficient neural networks from scratch by comprehensively exploring searchable units of multiple granularities and adaptively adjusting the retention ratios of the multi-granularity units. Additionally, we effectively tackle the challenge of high memory consumption of the search process without compromising the search quality via depth-level decomposition with progressive re-evaluation. Experimental results demonstrate the consistent superiority of MG-DARTS over other alternatives in terms of the accuracy-parameter trade-off. To further reduce the search cost, our future research efforts will focus on integrating zero-cost metrics to estimate the importance of searchable units at initialization.
VII Acknowledgments
This work was supported in part by the Hong Kong RGC General Research Fund (No.: PolyU 15205924); and in part by the Research Institute for Artificial Intelligence of Things, The Hong Kong Polytechnic University.